OpenStudy (anonymous):

Evaluate cot 45° without using a calculator by using ratios in a reference triangle.

OpenStudy (anonymous):

can someone please help i have no idea on what i have to do

jimthompson5910 (jim_thompson5910):

cot(x) = 1/tan(x) cot(45) = 1/(tan(45)) does that help?

OpenStudy (anonymous):

but how i do that .. i really didnt get this section

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

ok so the answer is goin to be tan(45)=opp/adj=1 ?

jimthompson5910 (jim_thompson5910):

that's tan(45)

jimthompson5910 (jim_thompson5910):

tan(45) = 1

jimthompson5910 (jim_thompson5910):

cot(x) = 1/tan(x) cot(45) = 1/(tan(45)) cot(45) = 1/(1) cot(45) = ???

OpenStudy (anonymous):

it 1 ?

jimthompson5910 (jim_thompson5910):

yep cot(45) = 1 also

OpenStudy (anonymous):

thank you so much :)

jimthompson5910 (jim_thompson5910):

np